Circular No. 8930

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2006 B7 (ODAS) Further to IAUC 8929, orbital elements for the comet’s last return: Epoch = 2005 Apr. 20.0 TT

T = 2005 May 3.1936 TT  ω = 69.^{o}1455
e = 0.447143  Ω = 358.6275 }2000.0
q = 1.980985 AU  i = 1.3516

a = 3.583176 AU n^{o} = 0.1453121 P = 6.78 years

S/2008 (41) 1 A. R. Conrad, W. M. Keck Observatory (WMKO); W. J. Merline,

Southwest Research Institute (SwRI); J. D. Drummond, Starfire Optical Range, Air Force Research Laboratory, Kirtland Air Force Base; P. M. Tamblyn, Binary Astronomy, Dillon, CO, and SwRI; C. Dumas, European Southern Observatory (ESO), Chile; B. X. Carry, ESO; R. D. Campbell and R. W. Goodrich, WMKO; W. M. Owen, Jet Propulsion Laboratory; and C. R. Chapman, SwRI, report the discovery on Mar. 28.5 UT of a satellite of minor planet (41) Daphne from more than 100 J-, H-, K-, and K′-band images with the 10-m Keck II telescope (+ NIRC2/AO adaptive-optics system) on Mauna Kea. On Mar. 28.5032, the satellite was at separation 0′′.56 (projected separation 443 km) and position angle 277^{o}. The satellite was observed to be moving with the primary, as they traveled 80′′ across the background sky over the 3 hours. The significant (0′′.3) orbital motion of the satellite observed during this time implies an orbital period near 1.6 days. The H-band brightness ratio is estimated at ∼ 10 mag, giving an estimated diameter of the satellite of < 2 km; if so, this system has the most extreme size ratio known (cf. EMP 2008, p. 256).

COMETS 196P/TICHÝ AND 197P/LINEAR Comet P/2008 C2 = 2000 U6 (cf. IAUC 8917) has been assigned the

number 196P (cf. MPC 61921), and comet P/2008 E2 = 2003 KV_{2} (cf. IAUCs 8924, 8925) has been assigned the number 197P (cf. MPC 62220).

STEREO SPACECRAFT Clarifying IAUC 8926, the two STEREO spacecraft will be 45^{o} from

the earth in its orbit about two years after launch, and continue to drift away from the earth (ultimately passing each other behind the sun).

2008 March 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green